OpenStudy (anonymous):

can someone graph this? xy = 15

OpenStudy (anonymous):

I would write it y = 15/x now pick some x values. at x=1 y= 15/1 = 15 at x= 15 y= 15/15 = 1 at x= 30 y = 15/30 = 1/2 at x= 1/2 y= 15/(1/2)= 30 so plot the points (1/2,30) (1,15) (15,1) (30, 1/2) after you plot the points connect the dots now try the same x numbers, but make them negative, and find the y numbers.